The average revenue function for a commodity is p = 50 – 4q. Find edp when:

demand = 5 units ; price = Rs. 6. Find consumer’s surplus at price = Rs. 6


1
Expert's answer
2021-05-12T12:19:45-0400

Solution:

AR function = Demand function

Edp = "\frac{P}{Q} \times \frac{\triangle Q}{\triangle P}"

Derive the inverse function of the demand function:

P = 50 – 4q

4q = 50 – P

q = 12.5 – "\frac{p}{4}"


"\frac{P}{Q} = \frac{6}{5}"


"\frac{\triangle Q}{\triangle P}" = -0.25


Edp = "(\frac{6}{5} ) \times -0.25 = -0.3"


Edp = -0.3

Consumer surplus = "\frac{1}{2}(Qd\times \triangle P)"

Derive Qd:

Price = 6

ΔP = Price the consumer is willing to pay – equilibrium price

Q = "12.5 - \frac{P}{4} = 12.5 - \frac{6}{4} = 12.5 - 1.5 = 11"

Qd = 11

The price the consumer is willing to pay = 50

Consumer surplus = "\frac{1}{2}(11\times (50 - 6)) = \frac{1}{2}(11\times 44) = \frac{1}{2}(484 ) = 242"


Consumer surplus = 242